Lois Bragg owns a small restaurant in Boston. Ms. Bragg provided her accountant with the following summary information regarding expectations for the month of June. The balance in accounts receivable as of May 31 is $57,000. Budgeted cash and credit sales for June are $145,000 and $586,000, respectively. Credit sales are made through Visa and MasterCard and are collected rapidly. Sixty percent of credit sales is collected in the month of sale, and the remainder is collected in the following month. Ms. Bragg's suppliers do not extend credit. Consequently, she pays suppliers on the last day of the month. Cash payments for June are expected to be $707,000. Ms. Bragg has a line of credit that enables the restaurant to borrow funds on demand; however, they must be borrowed on the last day of the month. Interest is paid in cash also on the last day of the month. Ms. Bragg desires to maintain a $34,000 cash balance before the interest payment. Her annual interest rate is 9 percent.

Required

  1. Compute the amount of funds Ms. Bragg needs to borrow for June.

  2. Determine the amount of interest expense the restaurant will report on the June pro forma income statement.

  3. What amount will the restaurant report as interest expense on the July pro forma income statement? (Round your answer to 2 decimal places.)

a. Amount to be borrowed
b. Interest expense (June)
c. Interest expense (July)
Answer #1

a)

Cash collection:
A Collection from May's receivables balance $57,000
B Collection from June's credit sales ($586,000 × 60%) $351,600
C Cash sales from June $145,000
D Total cash receipts (A+ B + C) $553,600
E Desired cash balance $34,000
F Cash Disbursement $707,000
G Total cash needs (E+F) $741,000
H Cash Shortage (G - D) $187,400

$187,400 is the amount of funds Ms. Bragg needs to borrow for June.

_______________________________________________________________________

b)

Interest expense (June) = $0, because amount is borrowed on the last day in the month of june.

______________________________________________________________________

c)

Interest expene (July) = Amount borrowed  \times 9% \times 1/12 months

= $187,400 \times 9% \times 1/12

= $1,405.5